OpenStudy (stephaun):

Solve for x: |2x + 6| − 4 = 20

OpenStudy (tkhunny):

Have you considered adding 4 and dividing by 2?

OpenStudy (stephaun):

Adding 4 to 20?

OpenStudy (zale101):

yes

OpenStudy (stephaun):

add 4 to 20 then divide 24 by 2

OpenStudy (zale101):

\[|2x+6|=24\] undue the absolute value term. To undue it you should follow this rule \[|x|=\pm x \] That means with your equation, you should use the plus or minus to the right side of the equation so u can undue the absolute value from the left side of the equation \[2x+6=\pm 24 \] Since 24 is plus or minus, then the equation will have two solutions when solving for x.

OpenStudy (zale101):

solve for x 2x+6=24 then solve for x when 24 is negative 2x+6=-(24)

OpenStudy (stephaun):

I got 2x=30 and 2x=-18

OpenStudy (zale101):

2x+6=24 subtract -6 to both sides u get 2x=18

OpenStudy (stephaun):

?

OpenStudy (zale101):

2x=30 is wrong too

OpenStudy (zale101):

2x+6=-(24) u subtract -6 to both sides 2x=-(24)-6 what's -24-6?

OpenStudy (stephaun):

-30

OpenStudy (zale101):

yes

OpenStudy (zale101):

for 2x+6=24 when u subtract -6 to both sides what will u get 2x+6(-6)=24-6 6-6 is zero so it cancels out 2x=24-6

OpenStudy (zale101):

24-6?

OpenStudy (stephaun):

So i would be 2x=-30 and 2x=18?

OpenStudy (zale101):

yep

OpenStudy (stephaun):

So it would be x=-15 and x=-9

OpenStudy (zale101):

now to isolate x for the multiplication of 2, divide 2 to both side 2x=-30 2/2x=-30/2 and 2x=18 2/2x=18/2

OpenStudy (zale101):

so 18/2=-9??

OpenStudy (stephaun):

Ohh ok i see now it would be x=-15 and x=9

OpenStudy (zale101):

yes

OpenStudy (stephaun):

ok thank you :3

OpenStudy (zale101):

yw